26%10:00 p.m × Question A worker at a fuel station knows from his experience that average nom of cars that arrive fir refiling at the station is 2 per minute. what is the probability that: a) there will be more than 2 cars in the next one minute. b) there will be atleat one car in the next 5 minutes
Answer #1

Whenever we have to estimate the probabilty of the number of cars arriving (arrival process), we model the situation using a Poisson distribution.

(a)

Let X denote the random variable representing the no. of cars arriving in a minute or arriving in the next one minute.

Now, we assume: X \sim Poi(\lambda)

\Rightarrow E(X) = \lambda

But we are given that the average number of cars arriving at the station in a minute is 2 cars.

Thus,

\begin{align*} E(X) &= 2 \\ \Rightarrow \ \ \ \ \ \lambda &= 2 \end{align*}

Thus, the probability mass function of X is given by:

P(X=x) = \frac{e^{-\lambda} \lambda ^x}{x!}=\frac{e^{-2}*2^x}{x!} \ \ \ ; x=0,1,2,...

Now, the required probability:

\begin{align*} \text{P(more than 2 cars in next minute)} &= P(X>2) \\ &= 1 - P(X \leq 2) \\ &= 1 - [P(X=0)+P(X=1)+P(X=2)] \\ &= 1 - [\frac{e^{-2}*2^0}{0!} + \frac{e^{-2}*2^1}{1!} + \frac{e^{-2}*2^2}{2!}] \\ &= 1 - [e^{-2} + 2e^{-2}+ 2e^{-2}] \\ &= 1-5e^{-2} \\ &= 1 - 5*0.135335 \\ &= 0.323324 \end{align*}which is the required answer.

(b)

Let Y denote the random variable representing the no. of cars arriving in 5 minutes or arriving in the next 5 minutes.

Now, we assume: Y \sim Poi(\mu)

\Rightarrow E(Y) = \mu

But we are given that the average number of cars arriving at the station in a minute is 2 cars.

Thus, the average number of cars arriving at the station in the next 5 minutes = 5*2 = 10 cars.

Thus,

\begin{align*} E(Y) &= 10 \\ \Rightarrow \ \ \ \ \ \mu &= 10 \end{align*}

Thus, the probability mass function of Y is given by:

P(Y=y) = \frac{e^{-\mu} \mu^y}{y!}=\frac{e^{-10}*10^y}{y!} \ \ \ ; y=0,1,2,...

Now, the required probability:

\begin{align*} \text{P(atleast one car in next 5 minutes)} &= P(Y \geq 1) \\ &= 1 - P(Y < 1) \\ &= 1 - P(Y=0) \\ &= 1 - \frac{e^{-10}*10^0}{0!} \\ &= 1- e^{-10} \\ &= 1 - 0.0000454 \\ &= 0.9999546 \end{align*}

which is the required answer.

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.